OpenStudy (anonymous):

find a quadratic equation whose roots are 3+i and -1/2

OpenStudy (precal):

can't do this because complex come in pairs 3+i and 3-i are solutions

OpenStudy (precal):

this is really a cubic

OpenStudy (anonymous):

right this que. is wrong

OpenStudy (precal):

were you asked to find a polynomial?

OpenStudy (precal):

if so then, we can do it

OpenStudy (anonymous):

no i was asked to find a quadratic. wouldn't the 3 + i and 3 - i pair only be required if i was asked for integral coefficients?

OpenStudy (anonymous):

If there is no requirement to have rational coefficients, then there's no problem, just multiply the factors together.

OpenStudy (anonymous):

*(meant to say "no requirement to have 'real' coefficients, ...")

OpenStudy (anonymous):

i know the formula x^2 - (sum of the roots)x + product of the roots = 0. but i did it and got x^2 - (5/2 + i)x - 2i = 0. the answer is supposed to be x^2-5+2i/2x-3+i/2

OpenStudy (anonymous):

@cliffsedge hello...?

OpenStudy (anonymous):

(Sorry, didn't see your reply) If the roots are 3+i and -1/2, the the factors are (x-3-i) and (x+1/2) When I multiply those together, I get \[\large x^2-\frac{5}{2}x-ix-\frac{3}{2}-\frac{1}{2}i\]

OpenStudy (anonymous):

I got that too. But the answer is supposed to be x^2-5+2i/2x-3+i/2

OpenStudy (anonymous):

I'm not sure I understand that answer. You mean this? \[\large x^2-\frac{(5+2i)}{2}x-\frac{(3+i)}{2}\] If so, then they are equivalent, try simplifying the above.

OpenStudy (anonymous):

no i meant this: